Circular No. 8123

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SAGITTARII 2003 AND V2377 SAGITTARII Independent discoveries on T-Max 400 films of an apparent nova have

been reported by Nicolas J. Brown (Quinns Rocks, W. Australia; mag 8.9 on Apr. 25.73 UT; α = 18^{h}40^{m}, δ = −33^{o}26′, equinox 2000.0) and by Minoru Yamamoto (Doi-machi, Okazaki, Japan, 200-mm f/4 camera lens; mag 9.6 on Apr. 26.735; α = 18^{h}40^{m}08^{s}.2, δ = −33^{o}26′37′′; via S. Nakano, Sumoto, Japan). Nothing was visible to mag 10.5 at this location on Yamamoto’s patrol films taken during 2000 Apr. 28–2003 Apr. 5 (though with limiting mag 9.5 on 2003 Mar. 11 and 29 films); the nearby Mira-type variable V2377 Sgr was not visible on any of these films. Nakano forwards the following accurate position for the new star from an unfiltered CCD exposure taken by R. Kushida (0.40-m reflector, Yatsugatake South Base Observatory; measured by Y. Kushida) on Apr. 27.677: α = 18^{h}40^{m}02^{s}.54, δ = −33^{o}26′55′′.1, mag 9.3; she adds that nothing is present at this position on Digitized Sky Survey (DSS) red and blue images. V2377 Sgr also appeared at mag 11.3 on Kushida’s exposure at α = 18^{h}40^{m}00^{s}.73, δ = −33^{o}27′49′′.6. Nakano also forwards position end figures 02^{s}.57, 26′55′′.5 (and mag 9.3) for the nova from an unfiltered CCD frame taken by K. Kadota (Ageo, Japan, 0.18-m reflector) on Apr. 27.688. A. C. Gilmore and P. M. Kilmartin, Mt. John University Observatory, report the following position end figures for N Sgr 2003 from a CCD exposure taken on Apr. 27 with the 0.6-m reflector: 02^{s}.55, 55′′.6 (uncertainty 0′′.3 in each coordinate). Kilmartin reports a possible precursor (apparently blended with a faint companion to its southeast) on a red DSS image at mag R = 17.9 ± 0.4 and position end figures 02^{s}.58, 55′′.3. Gilmore adds that L. Skuljan, University of Canterbury, obtained an échelle spectrogram of the new star with the Mt. John 1-m telescope around Apr. 27.67, the spectrum showing numerous broad and bright emission lines on a generally weak continuum. Photoelectric photometry of the nova by Gilmore with the 0.6-m reflector (reference stars Cousins E872 and E870): Apr. 27.667, V = 9.36 ± 0.01, U − B = −0.84 ± 0.03, B − V = −0.14 ± 0.03, V −R_{c} = +0.54 ± 0.01, V −I_{c} = +0.37 ± 0.02 (airmass 1.04). Visual magnitude estimates for the new nova: Apr. 26.67, 8.5 (Brown); 27.531, 9.0 (R. Stubbings, Drouin, Vic., Australia; via E. Waagen, AAVSO); 27.543, 9.3 (Gilmore); 27.738, 9.1 (A. Jones, Nelson, New Zealand).

SUPERNOVA 2001id IN UGC 12424 Corrigendum. On IAUC 7770, line 4, for 5′′.6 south read 5′′.6 north

2003 April 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green